### v1

# Estimating Bethe roots with VQE

Estimating Bethe roots with VQE
David Raveh111dxr921@miami.edu and  Rafael I. Nepomechie222nepomechie@miami.edu
Physics Department, PO Box 248046
University of Miami, Coral Gables, FL 33124 USA
Bethe equations, whose solutions determine exact eigenvalues and eigenstates of corresponding integrable Hamiltonians, are generally hard to solve. We implement a Variational Quantum Eigensolver (VQE) approach to estimating Bethe roots of the spin-1/2 XXZ quantum spin chain, by using Bethe states as trial states, and treating Bethe roots as variational parameters. In numerical simulations of systems of size up to 6, we obtain estimates for Bethe roots corresponding to both ground states and excited states with up to 5 down-spins, for both the closed and open XXZ chains. This approach is not limited to real Bethe roots.
1 Introduction
It is a remarkable fact that certain one-dimensional quantum many-body models, such as the spin-1/2 XXZ quantum spin chain, are quantum integrable, and can therefore be “exactly solved” by Bethe ansatz. (For reviews, see e.g. [1, 2, 3, 4].) The caveat is that the exact eigenvalues and eigenstates of such integrable Hamiltonians are expressed in terms of solutions (so-called Bethe roots) of a system of equations called Bethe equations; and the latter are generally hard to solve. Indeed, in addition to the venerable approach of Newton’s method, various classical methods have been developed to solve Bethe equations (see e.g. [5, 6, 7]); however, except when all the Bethe roots are real (in which case the Bethe equations can be efficiently solved by iteration), these methods generally work only for small system size $L$.
It is natural to ask whether quantum computers can provide an advantage for determining Bethe roots. A first attempt to address this question for the closed XXZ spin chain was made in [8], where it was suggested to use a Variational Quantum Eigensolver approach (see e.g. [9] and references therein) with Bethe states as trial states, treating Bethe roots as variational parameters. However, this effort did not go beyond $M=1$ (where $M$ is the number of down-spins), due to the difficulty of preparing the exact Bethe states. A probabilistic algorithm for preparing these Bethe states was then developed in [10], and it was subsequently generalized for the $U(1)$-invariant open XXZ chain in [11]. However, these algorithms were limited to real Bethe roots; and the success probability was shown to decrease super-exponentially with $M$ [12]. A deterministic algorithm for preparing Bethe states for the closed XXZ chain was developed in [13]; however, it required performing QR decompositions of matrices whose sizes scale exponentially with $M$. Analytic formulae for the unitaries of [13] were proposed in [14]. An explicit deterministic algorithm for preparing Bethe states for both the closed and open XXZ chains, for general values of $L$ and $M$, was recently formulated in [15], see also [16].
In the present work, we revisit the idea [8] of estimating Bethe roots by treating them as variational parameters in a VQE setup that uses Bethe states – prepared now with the new algorithm [15] – as trial states. In numerical simulations of systems of size up to $L=6$, we indeed obtain estimates for Bethe roots (some of which are complex) corresponding to both ground states and excited states with up to $M=5$ down-spins, for both the closed and open XXZ chain.
The outline of the remainder of the paper is as follows. In Sec. 2 we briefly review the closed and open XXZ chains and their corresponding Bethe equations, thereby also laying out our notations. In Sec. 3, we report results from our simulations for ground-state Bethe roots, obtained by minimizing the expectation value of the Hamiltonian using Bethe states as trial states. In Sec. 4, we report our results for excited-state Bethe roots, obtained by minimizing the Hamiltonian variance [17] using Bethe states as trial states. We end in Sec. 5 with a brief discussion of our results. Code in Qiskit [18] implementing these computations is available as Supplementary Material.
2 The models and their Bethe equations
We briefly review here the models and their corresponding Bethe equations, thereby also establishing our notations.
2.1 Closed chain
For the closed spin-1/2 XXZ quantum spin chain of length $L$ with periodic boundary conditions, we define the Hamiltonian
$\mathcal{H}=\tfrac{1}{4}\sum_{n=1}^{L}\left(\sigma^{x}_{n}\sigma^{x}_{n+1}+\sigma^{y}_{n}\sigma^{y}_{n+1}+\Delta\,\sigma^{z}_{n}\sigma^{z}_{n+1}\right)\,,\qquad\vec{\sigma}_{L+1}=\vec{\sigma}_{1}\,,$ (2.1)
where as usual $\sigma^{x}_{n},\sigma^{y}_{n},\sigma^{z}_{n}$ are Pauli matrices at site $n$, and $\Delta$ is the anisotropy parameter. The Hamiltonian has the $U(1)$ symmetry
$\left[\mathcal{H}\,,S^{z}\right]=0\,,\qquad S^{z}=\sum_{n=1}^{L}\tfrac{1}{2}\sigma^{z}_{n}\,.$ (2.2)
The corresponding Bethe equations for the so-called Bethe roots $\vec{k}=(k_{1},\ldots,k_{M})$ are given by (see e.g. [1, 2])  [^1] Bethe roots are often expressed instead in terms of variables $u_{j}$, which are related to $k_{j}$ by $e^{ik_{j}}=\frac{\sinh(u_{j}+\eta/2)}{\sinh(u_{j}-\eta/2)}$, where $\Delta=\cosh\eta$. The Bethe equations (2.3) then take the form $\left(\frac{\sinh(u_{j}+\frac{\eta}{2})}{\sinh(u_{j}-\frac{\eta}{2})}\right)^{L}=\prod_{l=1;l\neq j}^{M}\frac{\sinh(u_{j}-u_{l}+\eta)}{\sinh(u_{j}-u_{l}-\eta)}\,,\qquad j=1,\ldots,M\,.$
$e^{ik_{j}L}=\prod_{l=1;l\neq j}^{M}\left(-\frac{s(k_{l},k_{j})}{s(k_{j},k_{l})}\right)\,,\qquad j=1,\ldots,M\,,$ (2.3)
where
$s(k,k^{\prime})=1-2\Delta e^{ik^{\prime}}+e^{i(k+k^{\prime})}\,.$ (2.4)
The order of Bethe roots $k_{i}$ in $\vec{k}$ does not matter, since the Bethe equations are invariant under the interchange of any pair $(k_{j}\,,k_{l})$. Note also that the Bethe equations remain invariant under a $2\pi$ shift of any Bethe root, $k_{j}\mapsto k_{j}+2\pi$; hence, we can restrict $-\pi<\Re e(k_{j})\leq\pi$.
Indeed, the exact eigenstates of the Hamiltonian (2.1) are given (up to a normalization factor) by the Bethe states
$|B^{L}_{M}(\vec{k})\rangle=\sum_{w\in P(L,M)}f(w,\vec{k})\,|w\rangle\,,$ (2.5)
where the sum is over the set $P(L,M)$ of all permutations $w$ with $L-M$ zeros (up-spins) and $M$ ones (down-spins). Furthermore, $f(w,\vec{k})$ is given by
$f(w,\vec{k})=\sum_{\sigma\in S_{M}}\varepsilon(\sigma)\,A(k_{\sigma(1)},\cdots,k_{\sigma(M)})\,e^{i\sum_{j=1}^{M}k_{\sigma(j)}x_{j}}\,,$ (2.6)
where the sum is over the set $S_{M}$ of all permutations $\sigma$ of $\{1,\dots,M\}$, and $\varepsilon(\sigma)=\pm 1$ denotes the sign of the permutation $\sigma$. Moreover,
$A(k_{1},\ldots,k_{M})=\prod_{1\leq j<l\leq M}s(k_{l},k_{j})\,,$ (2.7)
and the $x_{j}\in\{1,\ldots,L\}$ in (2.6) are the positions of the 1’s in the argument $w$ of $f(w,\vec{k})$. The corresponding eigenvalues of $\mathcal{H}$ and $S^{z}$ are respectively given by
$E=\sum_{j=1}^{M}\cos(k_{j})+\left(\frac{L}{4}-M\right)\Delta\,,\qquad S^{z}=\frac{L}{2}-M\,.$ (2.8)
The Hamiltonian (2.1) has the charge conjugation (or duality) symmetry
$\mathcal{C}\,\mathcal{H}\,\mathcal{C}=\mathcal{H}\,,\qquad\mathcal{C}=\left(\sigma^{x}\right)^{\otimes L}\,,\qquad\mathcal{C}^{2}=1\,.$ (2.9)
The fact $\mathcal{C}\,S^{z}\,\mathcal{C}=-S^{z}$ implies a two-fold degeneracy for states with $S^{z}\neq 0$, and allows us to restrict $M\leq\lfloor L/2\rfloor$. [^2] If $|B^{L}_{M}(\vec{k})\rangle$ is an eigenstate with $M$ Bethe roots, then $\mathcal{C}\,|B^{L}_{M}(\vec{k})\rangle$ is an eigenstate with the same energy and opposite value of $S^{z}$, but with $L-M$ Bethe roots. (For $M=0$, the Bethe state is proportional to the reference state $|0\rangle^{\otimes L}$, where $|0\rangle=\binom{1}{0}$.)
2.2 Open chain
For the $U(1)$-invariant open spin-1/2 XXZ quantum spin chain of length $L$, we define the Hamiltonian
$\mathcal{H}=\tfrac{1}{4}\sum_{n=1}^{L-1}\left(\sigma^{x}_{n}\sigma^{x}_{n+1}+\sigma^{y}_{n}\sigma^{y}_{n+1}+\Delta\,\sigma^{z}_{n}\sigma^{z}_{n+1}\right)+\tfrac{1}{4}\left(h\,\sigma^{z}_{1}+h^{\prime}\,\sigma^{z}_{L}\right)\,,$ (2.10)
where again $\Delta$ is the anisotropy parameter, and $h$ and $h^{\prime}$ are boundary magnetic fields.
The corresponding Bethe equations for the Bethe roots $\vec{k}=(k_{1},\ldots,k_{M})$ are given by [19] [^3] In terms of the $u$-roots in Footnote 1, the Bethe equations (2.11) take the form $\frac{\sinh(u_{j}-\frac{\eta}{2}+a)}{\sinh(u_{j}+\frac{\eta}{2}-a)}\frac{\sinh(u_{j}-\frac{\eta}{2}+b)}{\sinh(u_{j}+\frac{\eta}{2}-b)}\left(\frac{\sinh(u_{j}+\frac{\eta}{2})}{\sinh(u_{j}-\frac{\eta}{2})}\right)^{2L}=\prod_{l=1;l\neq j}^{M}\left(\frac{\sinh(u_{j}-u_{l}+\eta)}{\sinh(u_{j}-u_{l}-\eta)}\frac{\sinh(u_{j}+u_{l}+\eta)}{\sinh(u_{j}+u_{l}-\eta)}\right)\,,$ where $h=\sinh\eta\coth b$ and $h^{\prime}=\sinh\eta\coth a$.
$\frac{\alpha(k_{j})\,\beta(k_{j})}{\alpha(-k_{j})\,\beta(-k_{j})}=\prod_{l=1;l\neq j}^{M}\frac{B(-k_{j},k_{l})}{B(k_{j},k_{l})}\,,\qquad j=1,\ldots,M\,,$ (2.11)
where
$B(k,k^{\prime})=s(k,k^{\prime})\,s(k^{\prime},-k)\,,$ (2.12)
see (2.4), and
$\alpha(k)=1+(h-\Delta)\,e^{-ik}\,,\qquad\beta(k)=\left[1+(h^{\prime}-\Delta)\,e^{-ik}\right]e^{i(L+1)k}\,.$ (2.13)
The Bethe equations (2.11) remain invariant under a $2\pi$ shift of any Bethe root, $k_{j}\mapsto k_{j}+2\pi$, as well as under the reflection $k_{j}\mapsto-k_{j}$ ; hence, we can restrict $0\leq\Re e(k_{j})\leq\pi$.
Indeed, the exact simultaneous eigenvectors of the Hamiltonian $\mathcal{H}$ (2.10) and $S^{z}$ again have the form $|B^{L}_{M}(\vec{k})\rangle$ (2.5), where $f(w,\vec{k})$ is now given by
$f(w,\vec{k})=\sum_{\sigma\in S_{M}}\sum_{\epsilon_{1},\dots,\epsilon_{M}=\pm 1}\varepsilon(\sigma)\,\epsilon_{1}\dots\epsilon_{M}\,A(\epsilon_{1}k_{\sigma(1)},\ldots,\epsilon_{M}k_{\sigma(M)})\,e^{i\sum_{j=1}^{M}\epsilon_{j}k_{\sigma(j)}x_{j}}\,.$ (2.14)
Moreover, now
$A(k_{1},\ldots,k_{M})=\prod_{j=1}^{M}\beta(-k_{j})\prod_{1\leq j<l\leq M}B(-k_{j},k_{l})\,e^{-ik_{l}}\,.$ (2.15)
The $x_{j}\in\{1,\ldots,L\}$ in (2.14) are again the positions of the 1’s in the argument $w$ of $f(w,\vec{k})$. The corresponding eigenvalues of $\mathcal{H}$ and $S^{z}$ are now given by
$E=\sum_{j=1}^{M}\cos(k_{j})+\left(\tfrac{1}{4}(L-1)-M\right)\Delta+\tfrac{1}{4}(h+h^{\prime})\,,\qquad S^{z}=\frac{L}{2}-M\,.$ (2.16)
For generic values of the parameters $\Delta$, $h$ and $h^{\prime}$, the spectrum is non-degenerate, and $M\in\{0,1,\ldots,L\}$. (For $M=L$, the Bethe state is proportional to the dual reference state $|1\rangle^{\otimes L}$, where $|1\rangle=\binom{0}{1}$.)
3 Ground-state Bethe roots
In this section, we apply the VQE algorithm, using Bethe states $|B^{L}_{M}(\vec{k})\rangle$ (2.5) as trial states and treating the Bethe roots $\vec{k}=(k_{1},\ldots,k_{M})$ as variational parameters, in order to estimate the Bethe roots corresponding to ground states. In this hybrid quantum/classical approach, where we aim to minimize the expectation value $\langle B^{L}_{M}(\vec{k})|\mathcal{H}|B^{L}_{M}(\vec{k})\rangle$, we evaluate the expectation values using a quantum simulator (ideally, a real device), using the algorithm in [15] to prepare the trial state $|B^{L}_{M}(\vec{k})\rangle$; and the variational parameters $\vec{k}$ are updated using a classical optimizer. We implement this approach using Qiskit v1.0 [18], with both the Statevector and Aer simulators, adopting cobyla as the classical optimizer.
We first consider in Sec. 3.1 the closed XXZ chain (2.1) with $L$ even in the antiferromagnetic regime $\Delta>0$, where the non-degenerate ground state is known to be described by $M=L/2$ real Bethe roots. We then consider in Sec. 3.2 the open XXZ chain (2.10), choosing values for the boundary fields such that the ground state is no longer described by only real Bethe roots.
3.1 Closed chain
We begin by implementing the VQE approach sketched above to the closed XXZ Hamiltonian (2.1) with $\Delta=2$, for even values of $L$ up to $L=6$. As already noted, the ground state is known to be described by $M=L/2$ real Bethe roots.
Our results are displayed in Table 1. The “Energy” is computed by direct diagonalization of the Hamiltonian (2.1); the “True roots” are computed by solving the Bethe equations (2.3) using Newton’s method, which is implemented in Mathematica by FindRoot; [^4] Since here the Bethe roots are real, they can be obtained more readily by solving the logarithm of the Bethe equations by iteration. the “Statevector roots” are computed using the Qiskit exact Statevector simulator; and the “Aer roots” are computed using the Qiskit noiseless shot-based Aer simulator with 10,000 shots.
We observe that the Statevector results are in excellent agreement with the true results. The Aer results are in fair agreement with the true results; increasing the number of shots improves the accuracy.
$L$ $M$ Energy True roots Statevector roots Aer roots
2 1 $-2$ $3.14159$ $3.1415$ $3.1487$
4 2 $-2.73205$$\pm 1.94553$ $1.9455,-1.9455$ $1.9623,-1.9503$
6 3 $-3.85577$$\pm 1.49862,3.14159$ $1.4986,-1.4986,3.1416$ $1.5477,-1.4830,3.1796$
Table 1: Ground-state Bethe roots for the closed chain (2.1)
3.2 Open chain
We now consider the open XXZ Hamiltonian (2.10) with $\Delta=1/2$, $h=3$, $h^{\prime}=3/10$, and values of $L$ up to $L=6$. The ground state is described here by $M=\lceil L/2\rceil$ Bethe roots; and the Bethe roots are not all real. Thus, for each complex root $k_{j}=\Re e(k_{j})+i\Im m(k_{j})$, two real variational parameters are introduced.
Our results are displayed in Table 2. The columns are computed similarly as in Table 1.
We observe that the Statevector results are again in excellent agreement with the true results; and the Aer results are again in fair agreement with the true results.
$L$ $M$ Energy True roots Statevector roots Aer roots
2 1 $-0.965015$ $3.14159+0.882174i$ $3.1415+0.8822i$ $3.0298+0.8785i$
3 2 $-1.49506$ $3.14159+0.908996i,$ $3.1412+0.9071i,$ $3.1537+0.9014i,$
$1.69883$ $1.6988$ $1.6930$
4 2 $-1.76803$ $3.14159+0.91503i,$ $3.1412+0.9142i,$ $3.0944+0.9929i,$
$2.11689$ $2.1169$ $2.1473$
5 3 $-2.22762$ $3.14159+0.916011i,$ $3.1399+0.9144i,$ $3.3120+1.5000i,$
$1.49569,2.31576$ $1.4958,2.3157$ $1.5449,2.3291$
6 3 $-2.53682$ $3.14159+0.916239i,$ $3.1419+0.9131i,$ $3.0864+0.9053i,$
$1.82675,2.47141$ $1.8266,2.4712$ $2.0142,2.3022$
Table 2: Ground-state Bethe roots for the open chain (2.10)
4 Excited-state Bethe roots
We turn now to the more challenging problem of estimating Bethe roots corresponding to excited states. To this end, we proceed similarly as before in Sec. 3, except rather than minimizing the expectation of the Hamiltonian $\langle B^{L}_{M}(\vec{k})|\mathcal{H}|B^{L}_{M}(\vec{k})\rangle$, we instead minimize the non-negative variance of the Hamiltonian
$\langle\left(\mathcal{H}-\langle\mathcal{H}\rangle\right)^{2}\rangle=\langle B^{L}_{M}(\vec{k})|\mathcal{H}^{2}|B^{L}_{M}(\vec{k})\rangle-\langle B^{L}_{M}(\vec{k})|\mathcal{H}|B^{L}_{M}(\vec{k})\rangle^{2}\,,$ (4.1)
which is 0 for any exact eigenstate, see e.g. [17]. Since a 0 variance can also arise if $|B^{L}_{M}(\vec{k})\rangle$ is null, we always check that $\langle B^{L}_{M}(\vec{k})|\mathcal{H}|B^{L}_{M}(\vec{k})\rangle$ is not 0.555This check fails for the exceptional cases that the energy is zero, such as the case $(L,M)=(2,1)$ in Table 3.
It should be noted that this method makes it equally likely to obtain the Bethe roots of any particular eigenstate; that is, there is no bias in favor of low-energy eigenstates. Thus, obtaining the Bethe roots of a particular eigenstate requires randomly selecting initial parameter values $\vec{k}$ until the output roots correspond to the desired eigenstate.
We first treat the closed chain in Sec. 4.1, followed by the open chain in Sec. 4.2.
4.1 Closed chain
We consider the closed XXZ Hamiltonian (2.1) again with $\Delta=2$, and values of $L$ up to $L=6$. Charge-conjugation invariance (2.9) allows us to restrict to $M\leq\lfloor L/2\rfloor$. Since complex roots of the closed-chain Bethe equations (2.3) always appear in conjugate pairs, we search for each such complex pair of Bethe roots using only two real variational parameters. As $L$ increases, the number of Bethe states rapidly increases; we therefore present only a representative sample.
Our results are displayed in Table 3. The columns are computed similarly as in Table 1. As for the ground-state Bethe roots, the Statevector results for excited-state Bethe roots are in excellent agreement with the true results, which confirms the viability of this variance-minimization approach. The Aer results are again in fair agreement with the true results.
We note that we do not find Bethe roots $\vec{k}$ that include $k=\pm i\infty$ (or $u=\mp\eta/2$ in terms of the $u$-roots in Footnote 1), which are called “physical singular solutions” in [6, 20]. Indeed, we do not expect to obtain such Bethe roots in this VQE approach, since a delicate limiting procedure is required to obtain the corresponding Bethe states, see e.g. [20] for the case $\Delta=1$.
$L$ $M$ Energy True roots Statevector roots Aer roots
2 1 00 00 00 $0.0019$
3 1 $-1$ $2.0944$ $2.0943$ $2.0938$
4 2 $0.732051$$\pm 0.831443i$$\pm 0.8314i$$\pm 0.8557i$
5 2 $0.716341$ $0.628319\pm 0.835459i$ $0.6276\pm 0.8349i$ $0.6263\pm 0.8858i$
6 2 $-1.75395$ $1.37766,2.81114$ $1.3776,2.8109$ $1.3998,2.8293$
6 3 $1.18614$ $0.244998\pm 1.41247i,$ $0.2451\pm 1.4120i,$ $0.2451\pm 1.3341i,$
$1.6044$ $1.6023$ $1.3152$
Table 3: Excited-state Bethe roots for the closed chain (2.1)
4.2 Open chain
We now consider the open XXZ Hamiltonian (2.10) again with $\Delta=1/2$, $h=3$, $h^{\prime}=3/10$, and values of $L$ up to $L=6$. Since this model does not have charge-conjugation invariance, we consider values of $M$ up to $M=L-1$. As for the closed chain, the number of Bethe states increases rapidly with $L$, so we present only a representative sample.
Our results are displayed in Table 4. The columns are computed similarly as in Table 1. Once again, we find excellent agreement between the Statevector and true results, and fair agreement between the Aer and true results.
Bethe roots $\vec{k}$ that include $k=\pi$ or $k=-i\infty$ (that is, $u=0$ or $u=\eta/2$, respectively, in terms of the $u$-roots in Footnote 1) do not appear in Bethe states of the open chain, see e.g. [21, 22]. Hence, in principle, it should be possible to obtain all open-chain Bethe roots by this VQE approach.
$L$ $M$ Energy True roots Statevector roots Aer roots
2 1 $0.715015$ $1.30258$ $1.3026$ $1.3025$
3 1 $-0.869852$ $3.14159+0.911371i$ $3.1417+0.9113i$ $3.1683+0.8783i$
4 2 $-0.224189$ $3.14159+0.916221i,$ $3.1413+0.9164i,$ $3.0311+0.9359i,$
$0.2264i$ $0.2284i$ $0.2148i$
4 3 $-0.128194$ $3.14159+0.916237i,$ $3.1416+0.9174i,$ $3.0528+0.9138i,$
$0.93789,0.245389i$ $0.9382,0.2474i$ $0.9046,0.2146i$
5 4 $-1.61607$ $3.14159+0.916185i,$ $3.1421+0.9027i,$ $3.0030+0.7578i,$
$0.514675,1.16211,2.43263$ $0.5044,1.1618,2.4332$ $0.4451,1.0290,2.4882$
6 5 $0.21968$ $3.14159+0.916291i,$ $3.1412+0.9156i,$ $3.1218+0.7523i,$
$0.667057,0.32195i$ $0.6252,0.3029i,$ $0.5055,0.5615i,$
$1.12044\pm 0.160175i$ $1.1347\pm 0.1956i$ $1.0101\pm 0.2531i$
Table 4: Excited-state Bethe roots for the open chain (2.10)
5 Discussion
We have seen in noiseless simulations that Bethe roots for both ground states and excited states of the XXZ chain can indeed be estimated using a VQE setup with Bethe states as trial states and Bethe roots as variational parameters. This approach is not limited to real Bethe roots, and can therefore be used to look for solutions of the Bethe equations that cannot be easily obtained classically by iteration.
A limitation of this approach, which is also present in the classical approach of solving Bethe equations using Newton’s method, is the need for the initial values of the Bethe roots to be “sufficiently close” to the true values. As the number of real variational parameters (which is $\mathcal{O}(M)$) increases, this problem is exacerbated. A further limitation is the presence of shot noise, which however can be ameliorated by increasing the number of shots. On the other hand, an advantage of our VQE approach for ground-state Bethe roots over directly solving Bethe equations is that, in the former approach, the resulting Bethe roots are guaranteed to be “physical”; that is, they correspond to actual eigenstates of the Hamiltonian. (When directly solving Bethe equations, one often finds “unphysical” solutions, such as those containing repeated roots, which do not correspond to actual eigenstates, see e.g. [6].)
The circuit [15] that we used to prepare the trial states renders both the classical and quantum parts of this VQE approach expensive. Indeed, for each VQE iteration, the computation of the $f(w,\vec{k})$ coefficients for the closed and open chains respectively require $M!$ and $2^{M}M!$ summations (see Eqs. (2.6) and (2.14)), making the classical part expensive for large $M$. Moreover, the circuit has size and depth $\mathcal{O}(\binom{L}{M}$) [15], making also the quantum part expensive for large $L$ and $M$. This VQE approach for estimating Bethe roots is therefore not feasible in the near term (the so-called Noisy Intermediate-Scale Quantum era). However, on fault-tolerant hardware, this approach may have an advantage over classical hardware for moderate values of $L$ and $M$.
It would be interesting to reduce the classical computational cost of the Bethe state circuit, perhaps by implementing a type of algebraic Bethe ansatz [13, 14], instead of the coordinate Bethe ansatz implemented in [15]. It would also be interesting to reduce the size and/or depth of the Bethe circuit, perhaps by exploiting ancillas, measurements, and feedforward operations as in [23] for Dicke states.
Acknowledgments
We thank John Van Dyke for valuable discussions, and Ananda Roy for reading the draft. We are also grateful to Daniel Nepomechie and Chana Raveh for their help with python. RN is supported in part by the National Science Foundation under grant PHY 2310594, and by a Cooper fellowship.